OpenStudy (anonymous):

what is the inverse of f if f(x)=^3 sqrt(x-5) answers: f^-1(x)=x^3+5 f^-1(x)=1/ ^3 sqrtx-5 f^-1(x)=x^3+125 f^-1(x)=5-^3 sqrtx

OpenStudy (anonymous):

to solve for the inverse you should flip your x and y and solve for x where f(x)=y

OpenStudy (anonymous):

but idont see a y

OpenStudy (anonymous):

\[y=\sqrt[3]{x-5}\] \[x=\sqrt[3]{y-5}\] there

OpenStudy (anonymous):

then \[x^3=y-5\\ x^3+5=y\] and that is your inverse

OpenStudy (anonymous):

so a?

OpenStudy (anonymous):

i stated where f(x)=y